OpenStudy (anonymous):

MEDALS FOR BEST ANSWERS. Find f(5) for this sequence f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) =

OpenStudy (ranga):

f(n) = f(1) + f(2) + f(n - 1), for n > 2 Substitute f(1) + f(2) into the formula first f(n) = 2 + 3 + f(n - 1) f(n) = 5 + f(n - 1) for n > 2 Put n = 3 f(3) = 5 + f(3 - 1) = 5 + f(2) = 5 + 3 = 8 f(4) = ? f(5) = ?

OpenStudy (anonymous):

find f(3)=f(1)+f(2)+f(3-1)=2+3+3=8 similarly we can put n =4 than n=5 to find f(5)

OpenStudy (anonymous):

so what will it be

OpenStudy (anonymous):

f(4)=f(1)+f(2)+f(4-1)=2+3+8=13 so f(5)=? do it like i found f(5)

OpenStudy (anonymous):

i don't get it at all that's why i had to ask

OpenStudy (anonymous):

ok but that should have been enough i will find it for u f(5)=f(1)+f(2) +f(4)=2+3+13=18

OpenStudy (anonymous):

thank you:)